The energy of motion is called: Kinetic energy.
Answer #1

1) a) Kinetic energy

2) a) Momentum is conserved

3) b) Doubled   [KE = 0.5*m*v^2]

4) c) Object is moving with constant speed           [W = F*d = ma*d]

5) c) Both KE and momentum are unchanged

6)c) Both a) and b)

7) c) Zero     [W = F*d*cos(theta)]

8) b) PE is being converted to KE

9) d) KE + PE conserved

10) b) net work done on the puck is -ve

11) b) square of velocity

12) c) 4K